Describe how energy requirements are determined or estimated. Try two different methods to calculate your energy needs (calories). Is there a difference between the two methods? Which method do you feel is most accurate? Explain your answer. Provide reference
Answer #1

Energy requirements can be calculated by considering the age,sex,weight ,height,and the range of physical activity he does.estimated energy requirements is the avarage calorie value which is needed for a person in a day.

One of the easiest methods to calculate energy need is depend on kilo calories per gram of body weight per day

The value will be constant

For normal individulal it is 25 to 30 kcal

Patient with small sickness 30 to 35 kcal.

Serious illness patient 35 to 40 kcal

My body weight - 65 kg

Constant value(normal person)-30 kcal/kg/day

Enery requirement in day=body weight(kcal/kg/day)

=65×30=1950 kcal/day

Other method is to find out age of person that is -31

For women 31 to 60 yrs weight multiply with 8.7 pl,us 829 (weight×8.7+829)

18-30 age( weight×14.7+496)

My age is 31 and weight 65 kg so the calculation is

(65×8.7)+829=1394.5

Reasoning metabolic rate=1394.5

This answer should be calculated with value of daily activity level.value of daily activity will be constant.

Sedentry life style-1.4

A little active-1.7

Verymuch active-2.0 these are constant values for everyone according to physical a tivity.

My life style is moderately active so

1394.5×1.7=2370.65

Total expenditure=2370.65

Average daily value=2370.65÷7

=338.6kcal

I think first method is easy and accurate